# Non-Hermitian Origin of Wannier Localizability and Detachable Topological Boundary States

### v2

@@ Abstract
Topological insulators and superconductors play a pivotal role in modern condensed matter physics [1, 2], generally classified by the fundamental tenfold internal symmetry classification [3, 4, *Ryu-10, 6, 7]. A distinctive feature of the topologically nontrivial bulk bands is the emergence of anomalous states at boundaries. Additionally, certain topology, such as the Chern number, imposes obstructions to constructing exponentially localized Wannier functions in the bulk bands [8, 9, 10, 11, 12]. The Wannier localizability also provides a foundation of topological phases, including topological crystalline insulators [13]. Conversely, other types of topology do not necessarily lead to the Wannier obstructions [14, 15, 16]. Remarkably, contrary to the conventional intuition of the bulk-boundary correspondence, such Wannier-localizable topological insulators have recently been shown to host boundary states that are detachable from the bulk bands [17, 18]. In this Letter, we reveal that these detachable topological boundary states originate from non-Hermitian topology.
@@
Table 1: Wannier localizability of the $d$-dimensional bulk and detachability of $\left(d-1\right)$-dimensional boundary states in topological insulators and superconductors. The tenfold Altland-Zirnbauer symmetry classes comprise time-reversal symmetry (TRS), particle-hole symmetry (PHS), and chiral symmetry (CS). The entries with $\checkmark$” (“$\times$”) accompany Wannier nonlocalizable (localizable) topological phases in the bulk and exhibit nondetachable (detachable) boundary states with intrinsic (extrinsic) non-Hermitian topology. The entries with “/$\checkmark$/$\times$ / ×” specify the Wannier nonlocalizable and localizable topological phases for the odd and even numbers of topological invariants, respectively.
@@
A 00 00 00 00 $\mathbb{Z}^{\checkmark}$ 00 $\mathbb{Z}^{\checkmark}$ 00 $\mathbb{Z}^{\checkmark}$ 00 $\mathbb{Z}^{\checkmark}$
AIII 00 00 $1$ $\mathbb{Z}^{\times}$ 00 $\mathbb{Z}^{\times}$ 00 $\mathbb{Z}^{\times}$ 00 $\mathbb{Z}^{\times}$ 00
AI $+1$ 00 00 00 00 00 $2\mathbb{Z}^{\checkmark}$ 00 $\mathbb{Z}_{2}^{\checkmark}$ $\mathbb{Z}_{2}^{\checkmark}$ $\mathbb{Z}^{\checkmark}$
BDI $+1$ $+1$ $1$ $\mathbb{Z}^{\times}$ 00 00 00 $2\mathbb{Z}^{\times}$ 00 $\mathbb{Z}_{2}^{\times}$ $\mathbb{Z}_{2}^{\times}$
D 00 $+1$ 00 $\mathbb{Z}_{2}^{\times}$ $\mathbb{Z}^{\checkmark}$ 00 00 00 $2\mathbb{Z}^{\checkmark}$ 00 $\mathbb{Z}_{2}^{\times}$
DIII $-1$ $+1$ $1$ $\mathbb{Z}_{2}^{\times}$ $\mathbb{Z}_{2}^{\checkmark}$ $\mathbb{Z}^{\checkmark/\times}$ 00 00 00 $2\mathbb{Z}^{\times}$ 00
AII $-1$ 00 00 00 $\mathbb{Z}_{2}^{\checkmark}$ $\mathbb{Z}_{2}^{\checkmark}$ $\mathbb{Z}^{\checkmark}$ 00 00 00 $2\mathbb{Z}^{\checkmark}$
CII $-1$ $-1$ $1$ $2\mathbb{Z}^{\times}$ 00 $\mathbb{Z}_{2}^{\times}$ $\mathbb{Z}_{2}^{\times}$$\mathbb{Z}^{\times}$ 00 00 00
C 00 $-1$ 00 00 $2\mathbb{Z}^{\checkmark}$ 00 $\mathbb{Z}_{2}^{\times}$ $\mathbb{Z}_{2}^{\times}$ $\mathbb{Z}^{\checkmark}$ 00 00
CI $+1$ $-1$ $1$ 00 00 $2\mathbb{Z}^{\times}$ 00 $\mathbb{Z}_{2}^{\times}$ $\mathbb{Z}_{2}^{\checkmark}$ $\mathbb{Z}^{\checkmark/\times}$ 00
Non-Hermitian topology.—A distinctive feature of non-Hermitian systems is complex-valued spectra, leading to point and line gaps [32, 41]. In the presence of a point (line) gap, complex-energy bands do not cross a reference point (line) in the complex-energy plane. Specifically, a non-Hermitian Hamiltonian $H$ is defined to have a point gap for $E_{n}\neq E_{\rm P}$, where $E_{n}$’s are complex eigenenergies of $H$, and $E_{\rm P}\in\mathbb{C}$ is a reference energy. By contrast, $H$ is defined to have a real (imaginary) line gap for $\mathrm{Re}\,E_{n}\neq E_{\rm L}$ ($\mathrm{Im}\,E_{n}\neq E_{\rm L}$) with $E_{\rm L}\in\mathbb{R}$, implying separable bands [28]. Even if a point gap is open, a line gap is not necessarily open. However, a point gap is always open when a line gap is open by placing a reference point on the reference line. Consequently, some point-gap topology is continuously deformable to line-gap topology and therefore has a counterpart in Hermitian or anti-Hermitian systems. In contrast to such extrinsic non-Hermitian topology, some point-gap topology is irreducible to line-gap topology and thus intrinsic to non-Hermitian systems.
@@
where $\gamma>0$ is the degree of non-Hermiticity and describes the nonreciprocal hopping in real space [88, *Hatano-Nelson-97]. This identification is formalized by adding dissipation to the edges [63, 64, 65]. The non-Hermitian term can also be understood as the coupling between the bulk and edges [66], where the bulk degrees of freedom at infinity correspond to decaying eigenstates with $\mathrm{Im}\,E<0$. Around the Fermi surface $\mathrm{Re}\,E=0$, this non-Hermitian system $H_{\rm A}\left(k_{x}\right)$ hosts two modes $k_{x}$ and $-k_{x}-2\text{i}\gamma$, only the former of which survives since the latter decays with a finite lifetime $1/2\gamma$ [48, 52].
@@
Specifically, $H_{\rm A}\left(k_{x}\right)$ in Eq. (1) is characterized by $W_{1}\left[H_{\rm A}\right]=1$ for the reference energy $E_{\rm P}$ inside the loop of the complex spectrum. Importantly, this complex-spectral winding always vanishes for Hermitian systems and is hence intrinsic to non-Hermitian systems. Thus, the point gap cannot be opened or reduced to any line gaps while preserving the point-gap topology. Consequently, the chiral edge state around $E=0$ cannot be detached from other eigenstates with $\mathrm{Im}\,E<0$ that play a role of the bulk degree of freedom. To further confirm this stability, we couple the chiral edge state to a trivial band $H=0$,
@@
with the coupling amplitude $v\in\mathbb{C}$. As momentum is traversed through the Brillouin zone, the perturbed complex two bands swap with each other, prohibiting the point-gap opening [Fig. 1 (b)].
Dirac surface state and extrinsic non-Hermitian topology.—In contrast to the chiral edge states, we demonstrate that the detachable topological boundary states accompany extrinsic non-Hermitian topology. As a prime example, we investigate a Dirac surface state of a three-dimensional topological insulator protected solely by chiral symmetry: $\mathcal{H}_{\rm bdy}\left(\bm{k}\right)=k_{x}\sigma_{x}+k_{y}\sigma_{y}$. This Dirac Hamiltonian indeed respects chiral symmetry $\Gamma\mathcal{H}_{\rm bdy}\left(\bm{k}\right)\Gamma^{-1}=-\mathcal{H}_{\rm bdy}\left(\bm{k}\right)$ with $\Gamma=\sigma_{z}$ and hence belongs to class AIII [7]. Since chiral symmetry is relevant only to zero energy, $\mathcal{H}_{\rm bdy}\left(\bm{k}\right)$ accompanies no spectral flow and is hence detachable from the bulk bands [18, 90].
@@
with $\Gamma=\sigma_{z}$. Around the Fermi surface $\mathrm{Re}\,E=0$ (i.e., $k_{x}=0$, $k_{y}=\pm\pi/2$), it comprises two modes, $k_{x}\sigma_{x}+k_{y}\sigma_{y}$ and $k_{x}\sigma_{x}-k_{y}\sigma_{y}-2\text{i}\gamma$, only the former of which survives and reduces to the Hermitian Dirac surface state $\mathcal{H}_{\rm bdy}\left(\bm{k}\right)$ [48, 52].
@@
A crucial distinction from the chiral edge states is that this point-gap topology is continuously deformable to imaginary-line-gap topology while preserving the point gap and chiral symmetry (Fig. 2). In fact, when we couple $H_{\rm AIII}\left(\bm{k}\right)$ to a trivial band $H=\varepsilon\sigma_{x}$ ($\varepsilon\geq 0$) by
@@
with $v\in\mathbb{C}$, the bands around $E=0$ are separated from the other bands with $\mathrm{Im}\,E<0$, leading to the opening of an imaginary line gap. As discussed above, decaying eigenstates $\mathrm{Im}\,E<0$ describe the bulk degrees of freedom. Thus, the imaginary-line-gap opening indicates the detachment of the Dirac surface state from the bulk bands.
Once the boundary states are detached, we can introduce a (non-Hermitian) Hamiltonian $H_{\rm bdy}\left(\bm{k}\right)$ projected onto the detached boundary bands [see, for example, the red dotted rectangle in Fig. 2 (c)], and define its imaginary-line-gap topology. Notably, chiral symmetry in Eq. (5) behaves like $\mathbb{Z}_{2}$ unitary symmetry that commutes with $H_{\rm bdy}\left(\bm{k}\right)$ [91]. Hence, $H_{\rm bdy}\left(\bm{k}\right)$ hosts two independent Chern numbers $\mathrm{Ch}_{1}\,[H_{\rm bdy}^{(\pm)}]$ in the subspaces of $\Gamma=\pm 1$. This $\mathbb{Z}\oplus\mathbb{Z}$ topological classification for imaginary line gaps includes the $\mathbb{Z}$ classification for point gaps, showing the extrinsic nature of the point-gap topology. Specifically, a map from $\mathbb{Z}\oplus\mathbb{Z}$ to $\mathbb{Z}$ is given as [51, 92, 90]
@@
where a reference point and line are chosen appropriately. Thus, in combination with Eq. (6), as a direct consequence of the bulk topology $W_{3}\left[\mathcal{H}_{\rm bulk}\right]\neq 0$, the detached boundary states inevitably exhibit the nontrivial Chern number $\mathrm{Ch}_{1}\,[H_{\rm bdy}^{(+)}]\neq 0$ or $\mathrm{Ch}_{1}\,[H_{\rm bdy}^{(-)}]\neq 0$, consistent with the surface Chern number introduced in Refs. [17, 18]. We summarize the connection between these topological invariants in Fig. 3. Notably, using this general discussion, we systematically classify potential nontrivial topology of detached boundary states (Table 2), as discussed below.
@@
Classification.—While we have hitherto investigated the prototypical models, our formulation is generally applicable to topological materials, leading to the classification of Wannier localizability and detachable topological boundary states in Table 1 (see also Fig. 4). First, we identify topological boundary states as non-Hermitian systems with point-gap topology [48, 52]. There, the Altland-Zirnbauer symmetry for original Hermitian systems reduces to the Altland-Zirnbauer symmetry for non-Hermitian systems [41], based on chiral symmetry in Eq. (5), as well as time-reversal symmetry and particle-hole symmetry,
@@
Then, the detachability of the topological boundary states reduces to whether the corresponding point-gap topology is intrinsic. This is classified based on homomorphisms from imaginary-line-gap topology to point-gap topology [51, 92, 93, 90]. According to this classification, all the topological boundary states in the Wigner-Dyson classes (i.e., classes A, AI, and AII) cannot be detached from the bulk bands, consistent with their spectral flow [94]. By contrast, in the chiral or Bogoliubov-de Gennes classes, some topological boundary states can be detached. This feature arises from extrinsic point-gap topology, further reducing to trivial phases in the Wigner-Dyson classes by ignoring chiral and particle-hole symmetries. Moreover, as a direct consequence of the bulk topology, the detached boundary states exhibit imaginary-line-gap topology, as summarized in Table 2. Similarly to the three-dimensional bulk in class AIII, the boundary topology is captured by $H_{\rm bdy}\left(\bm{k}\right)$, for which particle-hole symmetry behaves like time-reversal symmetry [i.e., $\mathcal{C}H_{\rm bdy}^{*}\left(\bm{k}\right)\mathcal{C}^{-1}=H_{\rm bdy}\left(-\bm{k}\right)$]. Notably, such a projected Hamiltonian can be introduced even for purely Hermitian detached boundary states [95, 90].
Our classification also predicts other possible topological phases that host detachable boundary states. For example, surface states in three-dimensional $\mathbb{Z}_{2}$ topological insulators for class CII are also detachable from the bulk bands. As an advantage of our formalism, topology of the detached boundary states is systematically identified based on the classification with respect to imaginary line gaps. In the projected subspace of the detached surface states in class CII, particle-hole symmetry in the original space behaves like time-reversal symmetry, leading to the $\mathbb{Z}_{2}$ Kane-Mele or Fu-Kane topological invariant (see also Table 2[96, *Kane-Mele-05b, 98, 99, 100, 7]. Similarly, in four dimensions, topological boundary states in the Wigner-Dyson classes cannot be separated from the bulk, while those in the chiral and Bogoliubov-de Gennes classes (i.e., classes CII and C) are detachable. Additionally, in classes DIII and CI, since particle-hole transformation flips chirality, the bulk winding number $W_{2n+1}\left[\mathcal{H}_{\rm bulk}\right]$ should be even if the boundary states are detached, resulting in the “$\mathbb{Z}^{\checkmark/\times}$ classification” in Table 1.
@@
Class $\left(d-1\right)$-dim. boundary top. inv. $=$ $d$-dim. bulk top. inv.
@@
Wannier localizability.—Whereas we have focused on the detachability of topological boundary states, we now clarify its connection with the bulk Wannier localizability, using $K$-theory [101] (see Fig. 4). For a gapped $d$-dimensional Hermitian system, we consider the bulk Hamiltonian $\mathcal{H}_{\rm bulk}\,({\bm{k}})$ with periodic boundary conditions and the boundary Hamiltonian ^$\hat{\mathcal{H}}_{\rm bdy}\,({\bm{k}}_{\parallel})$ with semi-infinite boundary conditions [95]. The corresponding $K$-groups underlying their topological classifications are related by the bulk-boundary correspondence [101, 102, 103, 104, 105, 106]. As explained previously, the boundary Hamiltonian ^$\hat{\mathcal{H}}_{\rm bdy}\,({\bm{k}}_{\parallel})$ exhibits no spectral flow and is thus detachable if and only if it is in trivial phases within the corresponding Wigner-Dyson classes by ignoring chiral and particle-hole symmetries. From a bulk perspective, by contrast, this reduction to the Wigner-Dyson classes dictates the presence or absence of exponentially localized bulk Wannier functions. In fact, if ^$\hat{\mathcal{H}}_{\rm bdy}\,({\bm{k}}_{\parallel})$ exhibits spectral flow on the boundary, $\mathcal{H}_{\rm bulk}\left({\bm{k}}\right)$ holds a nontrivial topological invariant even when chiral and particle-hole symmetries are ignored, which obstructs the exponentially localized Wannier functions. Conversely, the absence of boundary spectral flow permits exponential Wannier localization [18]. We further elaborate on the $K$-theory classification in our accompanying work [107].
@@
Discussions.—In this Letter, we elucidate that non-Hermitian topology underlies Wannier localizability and detachable boundary states in Hermitian topological materials, thereby establishing their classification (Table 1). Specifically, we demonstrate that while intrinsic non-Hermitian topology prohibits the detachment of topological boundary states, extrinsic non-Hermitian topology permits it and further leads to anti-Hermitian topology of the detached boundary states. Our formulation reveals that the detached topological boundary states concomitantly exhibit Wigner-Dyson-type gapped topology in total and chiral-type or Bogoliubov-de Gennes-type gapless topology at zero energy. They should exhibit unique Anderson localization and transitions, meriting further investigation. Additionally, our formulation is applicable to topological crystalline insulators. As a prime example, we investigate reflection-invariant second-order topological insulators [108, 109, 110, 111, 90]. It is also worthwhile to extend our formulation to detachable boundary states in other types of topological insulators, such as Hopf insulators [112, 17] and nonsymmorphic insulators [113, 114, *Shiozaki-16].

### v3

# Non-Hermitian Origin of Detachable Boundary States in Topological Insulators

@@ Abstract
Topological insulators and superconductors play a pivotal role in modern condensed matter physics [1, 2], generally classified by the fundamental tenfold internal symmetry classification [3, 4, *Ryu-10, 6, 7]. A distinctive feature of the topologically nontrivial bulk bands is the emergence of anomalous states at boundaries. Additionally, certain topology, such as the Chern number, imposes obstructions to constructing exponentially localized Wannier functions in the bulk bands [8, 9, 10, 11, 12]. The Wannier localizability also provides a foundation of topological phases, including topological crystalline insulators [13]. Conversely, other types of topology do not necessarily lead to the Wannier obstructions [14, 15, 16]. Remarkably, contrary to the conventional intuition of the bulk-boundary correspondence, such Wannier-localizable topological insulators have recently been shown to host boundary states that are detachable from the bulk bands [17, 18]. In this Letter, we reveal an intimate connection of these detachable topological boundary states with non-Hermitian topology.
@@
Table 1: Wannier localizability of the $d$-dimensional bulk and detachability of $\left(d-1\right)$-dimensional boundary states in topological insulators and superconductors. The tenfold Altland-Zirnbauer symmetry classes comprise time-reversal symmetry (TRS), particle-hole symmetry (PHS), and chiral symmetry (CS). The entries with $\checkmark$” ($\times$”) accompany Wannier nonlocalizable (localizable) topological phases in the bulk and exhibit nondetachable (detachable) boundary states with intrinsic (extrinsic) non-Hermitian topology. The entries with $\checkmark/\times$” specify the Wannier nonlocalizable and localizable topological phases for the odd and even numbers of topological invariants, respectively.
@@
A 0 0 0 0 $\mathbb{Z}^{\checkmark}$ 0 $\mathbb{Z}^{\checkmark}$ 0 $\mathbb{Z}^{\checkmark}$ 0 $\mathbb{Z}^{\checkmark}$
AIII 0 0 11 $\mathbb{Z}^{\times}$ 0 $\mathbb{Z}^{\times}$ 0 $\mathbb{Z}^{\times}$ 0 $\mathbb{Z}^{\times}$ 0
AI +1+1 0 0 0 0 0 $2\mathbb{Z}^{\checkmark}$ 0 $\mathbb{Z}_{2}^{\checkmark}$ $\mathbb{Z}_{2}^{\checkmark}$ $\mathbb{Z}^{\checkmark}$
BDI +1+1 +1+1 11 $\mathbb{Z}^{\times}$ 0 0 0 $2\mathbb{Z}^{\times}$ 0 $\mathbb{Z}_{2}^{\times}$ $\mathbb{Z}_{2}^{\times}$
D 0 +1+1 0 $\mathbb{Z}_{2}^{\times}$ $\mathbb{Z}^{\checkmark}$ 0 0 0 $2\mathbb{Z}^{\checkmark}$ 0 $\mathbb{Z}_{2}^{\times}$
DIII 1-1 +1+1 11 $\mathbb{Z}_{2}^{\times}$ $\mathbb{Z}_{2}^{\checkmark}$ $\mathbb{Z}^{\checkmark/\times}$ 0 0 0 $2\mathbb{Z}^{\times}$ 0
AII 1-1 0 0 0 $\mathbb{Z}_{2}^{\checkmark}$ $\mathbb{Z}_{2}^{\checkmark}$ $\mathbb{Z}^{\checkmark}$ 0 0 0 $2\mathbb{Z}^{\checkmark}$
CII 1-1 1-1 11 $2\mathbb{Z}^{\times}$ 0 $\mathbb{Z}_{2}^{\times}$ $\mathbb{Z}_{2}^{\times}$ $\mathbb{Z}^{\times}$ 0 0 0
C 0 1-1 0 0 $2\mathbb{Z}^{\checkmark}$ 0 $\mathbb{Z}_{2}^{\times}$ $\mathbb{Z}_{2}^{\times}$ $\mathbb{Z}^{\checkmark}$ 0 0
CI +1+1 1-1 11 0 0 $2\mathbb{Z}^{\times}$ 0 $\mathbb{Z}_{2}^{\times}$ $\mathbb{Z}_{2}^{\checkmark}$ $\mathbb{Z}^{\checkmark/\times}$ 0
Non-Hermitian topology.—A distinctive feature of non-Hermitian systems is complex-valued spectra, leading to point and line gaps [32, 41]. In the presence of a point (line) gap, complex-energy bands do not cross a reference point (line) in the complex-energy plane. Specifically, a non-Hermitian Hamiltonian $H$ is defined to have a point gap for $E_{n}\neq E_{\rm P}$, where $E_{n}$’s are complex eigenenergies of $H$, and $E_{\rm P}\in\mathbb{C}$ is a reference energy. By contrast, $H$ is defined to have a real (imaginary) line gap for $\mathrm{Re}\,E_{n}\neq E_{\rm L}$ ($\mathrm{Im}\,E_{n}\neq E_{\rm L}$) with $E_{\rm L}\in\mathbb{R}$, implying separable bands [28]. Even if a point gap is open, a line gap is not necessarily open. However, a point gap is always open when a line gap is open by placing a reference point on the reference line. Consequently, some point-gap topology is continuously deformable to line-gap topology and has a counterpart in Hermitian or anti-Hermitian systems. In contrast to such extrinsic non-Hermitian topology, certain point-gap topology is irreducible to line-gap topology and thus intrinsic to non-Hermitian systems, serving as the origin of the skin effect [50, 51].
@@
where $\gamma>0$ is the degree of non-Hermiticity and describes the nonreciprocal hopping in real space [88, *Hatano-Nelson-97]. Hereafter, we use $\mathcal{H}$ ($H$) for Hermitian (non-Hermitian) Hamiltonians. This identification is formalized by adding dissipation to the edges [63, 64, 65] or incorporating the coupling between the bulk and edges [66], where the bulk degrees of freedom at infinity correspond to decaying eigenstates with $\mathrm{Im}\,E<0$. Around the Fermi surface $\mathrm{Re}\,E=0$, this non-Hermitian system $H_{\rm A}\left(k_{x}\right)$ hosts two modes $k_{x}$ and $k_{x}-2\text{i}\gamma$, only the former of which survives since the latter decays with a finite lifetime $1/2\gamma$ [48, 52].
@@
Specifically, $H_{\rm A}\left(k_{x}\right)$ in Eq. (1) is characterized by $W_{1}\left[H_{\rm A}\right]=1$ for the reference energy $E_{\rm P}$ inside the loop of the complex spectrum. This complex-spectral winding always vanishes for Hermitian systems and is intrinsic to non-Hermitian systems. Thus, the point gap cannot be opened or reduced to any line gaps while preserving the point-gap topology. Consequently, the chiral edge state around $E=0$ cannot be detached from other eigenstates with $\mathrm{Im}\,E<0$ that play a role of the bulk degree of freedom. To further confirm this stability, we couple the chiral edge state to a trivial band $H=0$ [Fig. 1 (b)],
@@
with the coupling amplitude $v\in\mathbb{C}$. As momentum is traversed through the Brillouin zone, the two bands swap with each other, prohibiting the point-gap opening [90].
Dirac surface state and extrinsic non-Hermitian topology.—In contrast to the chiral edge states, we demonstrate that the detachable topological boundary states accompany extrinsic non-Hermitian topology. As a prime example, we investigate a Dirac surface state of a three-dimensional topological insulator protected solely by chiral symmetry: $\mathcal{H}_{\rm bdy}\left(\bm{k}\right)=k_{x}\sigma_{x}+k_{y}\sigma_{y}$. This Dirac Hamiltonian indeed respects chiral symmetry $\Gamma\mathcal{H}_{\rm bdy}\left(\bm{k}\right)\Gamma^{-1}=-\mathcal{H}_{\rm bdy}\left(\bm{k}\right)$ with $\Gamma=\sigma_{z}$ and hence belongs to class AIII [7]. Since chiral symmetry is relevant only to zero energy, $\mathcal{H}_{\rm bdy}\left(\bm{k}\right)$ accompanies no spectral flow and is detachable from the bulk bands [18, 90].
@@
with $\Gamma=\sigma_{z}$. Around the Fermi surface $\mathrm{Re}\,E=0$ (i.e., $k_{x}=0$, $k_{y}=\pm\pi/2$), it comprises two modes, $k_{x}\sigma_{x}+k_{y}\sigma_{y}$ and $k_{x}\sigma_{x}-k_{y}\sigma_{y}-2\text{i}\gamma$, only the former of which survives and reduces to $\mathcal{H}_{\rm bdy}\left(\bm{k}\right)$ [48, 52].
@@
A crucial distinction from the chiral edge states is that this point-gap topology is continuously deformable to imaginary-line-gap topology while preserving the point gap and chiral symmetry (Fig. 2). When we couple $H_{\rm AIII}\left(\bm{k}\right)$ to a trivial band $H=\varepsilon\sigma_{x}$ ($\varepsilon\geq 0$) by
@@
with $v\in\mathbb{C}$, the bands around $E=0$ are separated from the other bands with $\mathrm{Im}\,E<0$, leading to the opening of an imaginary line gap [90]. As discussed above, decaying eigenstates with $\mathrm{Im}\,E<0$ describe the bulk degrees of freedom. Thus, the imaginary-line-gap opening indicates the detachment of the Dirac surface state from the bulk bands.
Once the boundary states are detached, we can introduce a (non-Hermitian) Hamiltonian $H_{\rm bdy}\left(\bm{k}\right)$ projected onto the detached boundary bands [see the red dotted rectangle in Fig. 2 (c)], and define its imaginary-line-gap topology. Notably, chiral symmetry in Eq. (5) behaves like $\mathbb{Z}_{2}$ unitary symmetry that commutes with $H_{\rm bdy}\left(\bm{k}\right)$ [91]. Hence, $H_{\rm bdy}\left(\bm{k}\right)$ hosts two independent Chern numbers $\mathrm{Ch}_{1}\,[H_{\rm bdy}^{(\pm)}]$ in the subspaces of $\Gamma=\pm 1$. This $\mathbb{Z}\oplus\mathbb{Z}$ topological classification for imaginary line gaps includes the $\mathbb{Z}$ classification for point gaps, showing the extrinsic nature of the point-gap topology. A map from $\mathbb{Z}\oplus\mathbb{Z}$ to $\mathbb{Z}$ is given as [51, 92, 90]
@@
where a reference point and line are chosen appropriately. Thus, in combination with Eq. (6), as a direct consequence of the bulk topology $W_{3}\left[\mathcal{H}_{\rm bulk}\right]\neq 0$, the detached boundary states inevitably exhibit the nontrivial Chern number $\mathrm{Ch}_{1}\,[H_{\rm bdy}^{(+)}]\neq 0$ or $\mathrm{Ch}_{1}\,[H_{\rm bdy}^{(-)}]\neq 0$, consistent with the surface Chern number introduced in Refs. [17, 18]. We summarize the connection between these topological invariants in Fig. 3.
@@
Classification.—We extend our formulation to generic topological materials and classify Wannier localizability and detachable topological boundary states in Table 1 (see also Fig. 4). First, we identify topological boundary states as non-Hermitian systems with point-gap topology [48, 52]. The Altland-Zirnbauer symmetry for original Hermitian systems reduces to the Altland-Zirnbauer symmetry for non-Hermitian systems [41], based on chiral symmetry in Eq. (5), as well as time-reversal symmetry and particle-hole symmetry,
@@
Then, the detachability of the topological boundary states reduces to whether the corresponding point-gap topology is intrinsic. This is classified based on homomorphisms from imaginary-line-gap topology to point-gap topology [51, 92, 93, 90]. According to this classification, all the topological boundary states in the Wigner-Dyson classes (i.e., classes A, AI, and AII) cannot be detached from the bulk bands, consistent with their spectral flow [94]. By contrast, in the chiral or Bogoliubov-de Gennes classes, some topological boundary states can be detached. This feature arises from extrinsic point-gap topology, further reducing to trivial phases in the Wigner-Dyson classes by ignoring chiral and particle-hole symmetries. Moreover, as a direct consequence of the bulk topology, the detached boundary states exhibit imaginary-line-gap topology, as summarized in Table 2. Similarly to the three-dimensional bulk in class AIII, the boundary topology is captured by $H_{\rm bdy}\left(\bm{k}\right)$, for which particle-hole symmetry behaves like time-reversal symmetry [i.e., $\mathcal{C}H_{\rm bdy}^{*}\left(\bm{k}\right)\mathcal{C}^{-1}=H_{\rm bdy}\left(-\bm{k}\right)$]. Such a projected Hamiltonian can be introduced even for purely Hermitian detached boundary states [95, 90]. In contrast to Ref. [18], our formulation directly captures the detachability of topological boundary states and further facilitates the classification of detached boundary states.
Our classification also predicts other possible topological phases that host detachable boundary states. For example, surface states in three-dimensional $\mathbb{Z}_{2}$ topological insulators for class CII are also detachable from the bulk bands. As an advantage of our formalism, topology of the detached boundary states is systematically identified based on the classification with respect to imaginary line gaps. In the projected subspace of the detached surface states in class CII, particle-hole symmetry in the original space behaves like time-reversal symmetry, leading to the $\mathbb{Z}_{2}$ Kane-Mele or Fu-Kane topological invariant (see also Table 2[96, *Kane-Mele-05b, 98, 99, 100, 7]. Similarly, in four dimensions, topological boundary states in the Wigner-Dyson classes cannot be separated from the bulk, while those in the chiral and Bogoliubov-de Gennes classes (i.e., classes CII and C) are detachable. Additionally, in classes DIII and CI, since particle-hole transformation flips chirality, the bulk winding number $W_{2n+1}\left[\mathcal{H}_{\rm bulk}\right]$ should be even if the boundary states are detached, resulting in the $\mathbb{Z}^{\checkmark/\times}$ classification” in Table 1.
@@
Class $\left(d-1\right)$-dim. boundary top. inv. == $d$-dim. bulk top. inv.
@@
Wannier localizability.—Whereas we have focused on the detachability of topological boundary states, we now clarify its connection with the bulk Wannier localizability, using $K$-theory [101] (see Fig. 4). For a gapped $d$-dimensional Hermitian system, we consider the bulk Hamiltonian $\mathcal{H}_{\rm bulk}\,({\bm{k}})$ with periodic boundary conditions and the boundary Hamiltonian $\hat{\mathcal{H}}_{\rm bdy}\,({\bm{k}}_{\parallel})$ with semi-infinite boundary conditions [95]. The corresponding $K$-groups underlying their topological classifications are related by the bulk-boundary correspondence [101, 102, 103, 104, 105, 106]. As explained previously, the boundary Hamiltonian $\hat{\mathcal{H}}_{\rm bdy}\,({\bm{k}}_{\parallel})$ exhibits no spectral flow and is thus detachable if and only if it is in trivial phases within the corresponding Wigner-Dyson classes by ignoring chiral and particle-hole symmetries. From a bulk perspective, by contrast, this reduction to the Wigner-Dyson classes dictates the presence or absence of exponentially localized bulk Wannier functions. In fact, if $\hat{\mathcal{H}}_{\rm bdy}\,({\bm{k}}_{\parallel})$ exhibits spectral flow on the boundary, $\mathcal{H}_{\rm bulk}\left({\bm{k}}\right)$ holds a nontrivial topological invariant even when chiral and particle-hole symmetries are ignored, which obstructs the exponentially localized Wannier functions. Conversely, the absence of boundary spectral flow permits exponential Wannier localization [18]. We further elaborate on the $K$-theory classification in our accompanying work [107].
@@
Discussions.—In this Letter, we elucidate that non-Hermitian topology underlies Wannier localizability and detachable boundary states in Hermitian topological materials, thereby establishing their classification (Table 1). We demonstrate that while intrinsic non-Hermitian topology prohibits the detachment of topological boundary states, extrinsic non-Hermitian topology permits it and further leads to anti-Hermitian topology of the detached boundary states. Our formulation reveals that the detached topological boundary states concomitantly exhibit Wigner-Dyson-type gapped topology in total and chiral-type or Bogoliubov-de Gennes-type gapless topology at zero energy. They should exhibit unique Anderson localization and transitions, meriting further investigation. Our framework may also be useful in engineering topological flat bands at boundaries. The precise conditions to realize boundary topological flat bands are left for future work.
Our non-Hermitian perspective directly characterizes the detachability of topological boundary states and complements the bulk classification of Wannier localizability [18]. This complementary viewpoint should be relevant to extending our results to topological crystalline insulators, where the bulk-boundary correspondence becomes more subtle. As a prime example, we investigate reflection-invariant second-order topological insulators [108, 109, 110, 111, 90]. It is also worthwhile to generalize our formulation to detachable boundary states in other types of topological insulators, such as Hopf insulators [112, 17] and nonsymmorphic insulators [113, 114, *Shiozaki-16]. Finally, a possible connection with topological unitary operators [116] is worth noting.